OpenStudy (anonymous):

Calculate the area of the circle whose diameter is 6 km. Note: you will need the radius. Please round to the nearest square kilometer. my answer was first to find the radius we divide the diameter in half 6/2= 3 so the radius is 3 so now we times that by x 32 = then i simlifly it to 9 then i subsitute 3.14 for 9 x 3.14 = 28.2 the area is exactly 9 or about 28.8 centimeters but my teacher says i need to rounded and the units should be squared

OpenStudy (amistre64):

pi (3)^2 is right for the area of the circle

OpenStudy (anonymous):

the radius of a circle is 1/2 the diameter, so the radius of this circle is 3. The formula for the area of a circle is: A = πr^2 where r is the radius approximating π as 3.14, Area = 9(3.14) = 28.26 sq. km. I do not know what you mean by "round to the n".

OpenStudy (amistre64):

the area is exactly 9pi; the approximation is then 28.2

OpenStudy (amistre64):

oh, units are what you used to measure with :) centimeters ... and area is always ^2

OpenStudy (amistre64):

ugh, km not cm

OpenStudy (amistre64):

... byt the time i got to readin gto the end i had forgotten what it started with ;)

OpenStudy (anonymous):

i think @amistre mean't this The diameter is related to the radius by, r = d/2 Therefore, our radius is, r = 6/2 = 3km The area of a circle is then, A = pi*r^2 = pi*(3^2) = 28.3 km^2

OpenStudy (anonymous):

thanks you guys are the best :)

OpenStudy (anonymous):

Welcome! :)